Phone was thrown from top of a building. The height h(t) of the cell phone at time (t) is described by the function h(t) = -16tsquared +96t+112. When will the phone be at a height of 192 feet and how many seconds will it take to hit the ground?
I have tried putting the numbers in and just cannot figure this out. Can you help please? Thanks

Answer by ankor@dixie-net.com(15652)

Phone was thrown from top of a building.
The height h(t) of the cell phone at time (t) is described by the function h(t) = -16tsquared +96t+112.
When will the phone be at a height of 192 feet and how many seconds will it take to hit the ground?
:
-16t^2 + 96t + 112 = h(t)
The phone will be at 192 ft when h(t) = 192
-16t^2 + 96t + 112 = 192
Subtract 192 from both sides
-16t^2 + 96t + 112 - 192
-16t^2 + 96t - 80 = 0; a quadratic equation
Simplify, and change the signs, divide each term by -16
t^2 - 6t + 5 = 0
Factors to
(t - 5)(t + 1) = 0
The positive solution
t = 5 sec , tel at 192 ft
:
Phone will hit the ground when h(t) = 0
-16t^2 + 96t + 112 = 0
Simplify again, divide by -16
t^2 - 6t - 7 = 0
Factors to
(t - 7)(t + 1) = 0
positive solution
t = 7 sec, phone strikes the ground